Tank A initially contains 200, litres of brine containing 225 N of salt. Eight litres of fresh water per A and the mixture, assumed uniform, passes from A to b. initially .containing 200 litres of fresh water, at 8 litres per minute. The resulting mixture, also kept uniform, leaves B at the rate of 8 litres/min. Find the amount of minute enter salt in tank B after one hour.


1
Expert's answer
2021-11-29T14:53:29-0500

Let "A(t) =" amount, in N of salt in tank A at time "t." Then we have

"\dfrac{dA}{dt}="(rate of salt into tank A) − (rate of salt out of tank A)


"\dfrac{dA}{dt}=0-\dfrac{8A}{200}"

So we get the differential equation


"\dfrac{dA}{A}=-\dfrac{dt}{25}"

Integrate

"A(t)=Ce^{-t\/25}"

Given "A(0)=225." Then


"A(t)=225e^{-t\/25}"

Let "B(t) =" amount, in N of salt in tank B at time "t." Then we have

"\dfrac{dB}{dt}="(rate of salt into tank B) − (rate of salt out of tank B)

"\dfrac{dB}{dt}=\dfrac{8}{200}(225e^{-t/25})-\dfrac{8}{200}B"

"\dfrac{dB}{dt}+0.04B=e^{-0.04t}"

"\dfrac{dB}{dt}+0.04B=0"


"\dfrac{dB}{B}=-0.04dt"

"B=c_1e^{-0.04t}"

"\dfrac{dB}{dt}=\dfrac{dc_1}{dt}e^{-0.04t}-0.04c_1e^{-0.04t}"

"\dfrac{dc_1}{dt}e^{-0.04t}-0.04c_1e^{-0.04t}+0.04c_1e^{-0.04t}"

"=e^{-0.04t}"

"\dfrac{dc_1}{dt}=1"

"c_1=t+c_2"


"B(t)=(t+c_2)e^{-0.04t}"

Given "B(0)=0." Then


"B(t)=te^{-0.04t}"

Hence


"B(60)=60e^{-0.04(60)}\approx5.443 \ N"